40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 21 02:42:17 2014, rana, ericq, Configuration, Green Locking, Gain changes on Green Y PDH TEK00002.PNGTEK00003.PNG
    Reply  Thu Aug 21 19:04:52 2014, ericq, Configuration, Green Locking, Gain changes on Green Y PDH TEK00005.PNG
       Reply  Tue Aug 26 23:16:49 2014, ericq, Configuration, Green Locking, Gain changes on Green Y PDH Ytfs.pdfpdhComparison.pdfnewY.zip
Message ID: 10418     Entry time: Thu Aug 21 02:42:17 2014     Reply to this: 10420
Author: rana, ericq 
Type: Configuration 
Category: Green Locking 
Subject: Gain changes on Green Y PDH 

[rana, ericq]

We spent time trying to relieve the Yend green PDH of it troubles. 

We realized that the mixer in the PDH setup (mini circuits ZAD-8+), wants 7dBm of LO to properly function. However, we use one function generators output, through a splitter, to give signals to the laser PZT and the mixer LO. 

We don't want 7dBm of power hitting the laser PZT, though. The summing node that adds the servo output to the sideband signal was supposedly designed to do some of this attenuation. Rana measured that 10Vpp out of the function generator resulted in 20mVpp on the fast input to the NPRO, after the summing node. Hence, the 0.09V setting was only resulting in something like 0.2mV hitting the PZT. The PZT has something like 30 rad/V PM response, meaning we only had ~0.006 rad of modulation. 

Now, the function generator is set to 2 Vpp, meaning 4 mVpp hitting the PZT, meaning ~0.12 radians of modulation. The mixer is now getting +7dBm on its LO, and the PDH traces look much cleaner. However, the PDH error signal is now something like 100mVpp, which is much bigger than the PDH board is designed for, so there is now a 10dB attenuator between the reflection PD DC block and the RF input to the mixer. 

Here are screenshots of the Inmon channel (which has a gain of ~20) showing a sweep through some PDH signal, and the error signal while in green lock. Huge 60Hz harmonics are still observed. 

TEK00002.PNGTEK00003.PNG

 


Regarding these 60Hz issues, we need to make sure that we remove all situations where long BNCs are chained together with barrel connectors, or Ts are touching other ones. We also should glue or affix the pomona summing box to the shelf, so that its not just laying on the floor.

The concrete next step is to go fiddle with things, and see if we can get the 60Hz noise to go away, then measure the PDH loop and noises again. Hopefully, this should make the ALS much more reliable. 

ELOG V3.1.3-